Please adhere to the rounding instructions(in red) for calculations and final answer. If the answer is not exact, it will be incorrect. Thanks!

D Question 14 1 pts You are planning to save for retirement over the next 33 years. To do this, you will invest $548 per month in a stock account and $252 per month in a separate bond account. The return of the stock account is expected to be 12%, and the bond account will pay 6%, when you retire, you will combine your money into an account with an expected 9% return. How much can you withdraw each month in retirement from your account assuming a 20-year withdrawal period? (Round answer to 2 decimal places. Do not round intermediate calculations Topic: Discounted Cash Flow Valuation
Answer #1
FV of the investment in the stock account = 548*(1.01^396-1)/0.01 = $   27,63,872.23
FV of the investment in the bond account = 252*(1.005^396-1)/0.005 = $     3,12,837.76
Total value of the investment at the end of 33 years = $   30,76,709.99
Amount that can be withdrawn each month for 20 years = 3076709.99*0.0075*1.0075^240/(1.0075^240-1) = $         27,681.96